# LOCUS – LOcalization with Channel Uncertainty and Sporadic Energy

### v2

@@ 4.1. Information-Weighted Focus (InFo)
Figure 7. Step-by-step workflow of Guided Replacement (GRep) to estimate the full-rank feature $\hat{F}$ from the low-rank feature $\tilde{F}$.
The first component of InFo is the channel information branch, $f_{\gamma}(\tilde{F})$, parameterized with $\gamma$. This branch performs average pooling followed by a multi-layer perceptron to extract the channel information entropy. Average pooling aggregates information across all indices, enabling the model to learn more robust features while accurately depicting the overall strength of a feature. The multi-layer perceptron consists of three fully connected layers with batch normalization and rectified linear unit (ReLU) activation after the first two layers. This branch outputs a vector $I_{c}=f_{\gamma}(\tilde{F})$ that measures global information for each channel.
The second component of InFo is the point-wise information branch, $f_{\lambda}(\tilde{F})$, parameterized with $\lambda$. The input $\tilde{F}$ first passes through a convolution layer with a kernel size of $(1\times 1)$ and a channel reduction factor of $r$. It then goes through $n$ dilated convolution layers with a kernel size of $(3\times 3)$, dilation value $d$, batch normalization, and ReLU activation. In our architecture, $n=3$ and $d=2$. The dilated convolution increases the receptive field, enabling us to leverage contextual information. Finally, this contextual information passes through another convolutional layer with a kernel size of $(1\times 1)$ and outputs $I_{s}=f_{\lambda}(\tilde{F})$, which estimates the point-wise information of the input feature.
Next, we broadcast $I_{s}$ to $\mathbb{R}^{d_{\tilde{X}}}$ to match the dimension of $\tilde{F}$ and perform element-wise addition on $I_{c}$ and $I_{s}$ to estimate the information entropy. Finally, the sigmoid function, $\sigma$, maps the results between 0 and 1, where a value closer to 1 indicates higher information availability and vice versa. Therefore, the information entropy is defined as:
@@
4.2. Latent Feature Synthesizer (LaFS)
@@
4.3. Guided Replacement (GRep)
The Guided Replacement (GRep) module recovers the full-rank feature $\hat{F}$ using Equation 3. Figure 7a depicts this information retrieval process. First, we perform element-wise multiplication between the information entropy $I(\tilde{F})$ and the low-rank feature set $\tilde{F}$ to retrieve the available information in the low-rank feature set $\hat{F}_{\tilde{X}}$.
Next, we find the missing information matrix $\hat{F}_{Z|\tilde{X}}$ from the interpolated feature $\bar{\bar{F}}$. Here, $Z$ represents the missing data in $\tilde{S}$. As shown in Figure 7b, we take the complement of the element-wise information matrix $I(\tilde{F})$ and multiply it element-wise with $\bar{\bar{F}}$ to find the missing information, disregarding regions (dotted darker green) where errors might have been introduced during interpolation.
@@ 4.3. Guided Replacement (GRep)
5. Experimental Setup and Implementation Details
This section details the dataset, data preparation, implementation, baselines, and evaluation metrics for assessing LOCUS. Our implementation can be found here (Hidethyself, [n. d.]).
@@ 5. Experimental Setup and Implementation Details
{tblr}
cells = c, row1 = Mercury, row3 = Mercury, Dataset & Duration $F_{s}$(Hz) No of samples Train-Val-Test
@@ 5.1. Dataset
To evaluate LOCUS, we utilized two distinct multichannel sound source localization datasets, each offering unique challenges and characteristics. These datasets are summarized in Table 1 and detailed below.
Dataset 1: DCASE. The first dataset is DCASE2021 Task 4 (Politis et al., 2021), comprising 600 four-channel, one-minute spatial recordings at 24 kHz. This polyphonic dataset includes up to four simultaneous sound sources, 12 different acoustic events, and provides Direction of Arrival (DoA) for each source. It also features directional interference and multi-channel ambient noise, with SNR levels ranging from 6 dB to 30 dB. The dataset is split into 400 development and 200 evaluation samples.
Dataset 2: LargeSet. While large single-channel acoustic datasets like Google Audioset (Gemmeke et al., 2017) with over 2 million annotated audio clips exist for sound event detection, they do not suffice for multichannel sound source localization. To address this gap, we created and published a 50-hour, six-channel audio dataset (Authors, 2023). This dataset includes 180,000 one-second recordings at 44.1 kHz across 10 different environments (Table 2). It is monophonic, ensuring no overlapping sound sources to isolate the impact of missing information.
Simulation Details. Using the PyRoomacoustics (Scheibler et al., 2018), we position a spherical microphone array (4.25 cm radius, six microphones) at the room’s geometric center to capture realistic spatial audio data. Sound source locations are randomly selected to simulate acoustic features, including room impulse responses (RIR) and noise. Different materials emulate diverse RIRs, with RT60 measuring the decay of reverberated sound. The UrbanSound8K dataset (Salamon et al., 2014) introduces real-world noise, with the signal-to-noise ratio (SNR) randomly varied to maintain realism. Further details are available in (Authors, 2023).
@@
{tblr}
colspec=X[0]X, rowodd = Mercury, Dimension & Description
@@ 5.2. Data Preparation
Feature Extraction. In our approach to sound source localization, we utilize two key acoustic features: Mel-Frequency Cepstral Coefficients (MFCC) and Generalized Cross Correlation (GCC). We extract these features from the raw audio data using a 1024-point Fast Fourier Transform (FFT) with a 40 ms window length and a 20 ms hop length at sampling rates of 24 kHz and 44.1 kHz. MFCC, calculated from 64 Mel bands, captures magnitude-based audio information, while GCC represents the relative time difference and phase between audio channels, crucial for determining the spatial origin of sound. We chose these features because MFCC provides important spectral details for distinguishing sound sources, while GCC captures spatial timing differences, enabling accurate localization. After truncating the GCC sequences to match the length of the Mel bins, we combine the two features into a feature matrix, $\tilde{F}$, which serves as input to deep learning pipeline for robust sound source localization. Integrating MFCC and GCC features enhances the network’s ability to accurately localize sound sources in complex acoustic environments.
Missing Channel Perturbation To prepare data with missing microphone channels, we randomly select $m\leq\frac{n}{2}$ channels from a total of $n$ microphones, where $m$ represents the number of malfunctioning or missing channels. For each Missing Data Percentage (MDP), we replace $p\%$ of the original audio with Gaussian noise, characterized by a mean of 0 and a standard deviation of 1. This noise simulates signal distortion or degraded microphone output. In addition, we simulate channel failures by randomly dropping entire channels for certain time frames (100% MDP), effectively mimicking intermittent outages where all microphones temporarily stop recording.
These perturbation techniques, including the introduction of Gaussian noise and random channel drops, ensure the model can handle real-world scenarios where microphone channels may fail or capture distorted signals.
@@ 5.3. Implementation Details
Training Pipeline of LOCUS. We train LOCUS jointly with the DNN-based sound source localization (SSL) task. The loss function for the SSL task, denoted as $\mathcal{L}_{DT}$, backpropagates through the SSL network and the channel and point-wise information branches ($f_{\gamma}$ and $f_{\lambda}$) of InFo. This ensures that the SSL task benefits from the enriched features provided by these branches.
However, the LaFS branch ($f_{\phi}$) has a separate training objective which is to estimate the input features. The loss function for DCI, $\mathcal{L}_{DCI}$, is defined as the squared $L_{2}$ distance between the original feature matrix $F$ and the reconstructed feature matrix $\bar{\bar{F}}$:
(5) $\mathcal{L}_{DCI}=||F-\bar{\bar{F}}||^{2}$
This loss function ensures that the reconstructed features closely match the original features, promoting consistency and robustness in the data representation.
To optimize the training process, we use two different optimizers: one for training InFo and the SSL task with $\mathcal{L}_{DT}$, and another for training the LaFS branch with $\mathcal{L}_{DCI}$. This dual-optimizer approach allows us to fine-tune each component of the system independently, ensuring that both the SSL task and the LaFS branch achieve optimal performance.
Sound Source Localization Pipeline.
We implement two DNN-based SLL solutions to estimate the Direction of Arrival (DoA) of sound sources using multi-channel audio – SELDNet (Shimada et al., 2021; Adavanne et al., 2018) and SALSA (Nguyen et al., 2022). SELDNet and SALSA both use MFCC as described in Section 5.2.
@@
SALSA is an advanced SSL algorithm featuring a CRNN model with a CNN backbone based on the PANN ResNet22 (Kong et al., 2020) architecture for audio tagging. It includes a two-layer Bidirectional GRU and fully connected output layers. The algorithm adapts to various input features by adjusting the number of input channels in the first convolutional layer. During inference, sound classes with probabilities above the detection threshold are considered active, and their corresponding DOAs are selected.
@@ 5.4. Evaluation Baseline
We rigorously evaluate LOCUS by benchmarking it against four established imputation methods and one deep interpolation technique. The selected baselines encompass a spectrum of strategies, from straightforward imputation techniques to sophisticated neural network models, ensuring a thorough and robust comparison. This comprehensive evaluation framework allows us to precisely assess the effectiveness and adaptability of LOCUS in handling missing data scenarios, highlighting its strengths and potential advantages over existing solutions.
@@
Imputation Methods. We benchmark LOCUS against three popular signal processing and two state-of-the-art deep learning imputation methods, one of which is specifically designed for acoustic signals. Each method is selected for its unique approach and effectiveness in handling missing data. Additionally, we evaluate each method on both pre-trained and retrained SSL networks.
@@
$\bullet$ Mean Imputation ($Mean$). Missing channel information is replaced with the mean of the available channels’ information (Allison, 2001). This method is chosen for its simplicity and widespread use as a baseline in missing data problems.
@@
$\bullet$ Hot Deck Imputation ($HotDeck$). This method fills in missing channel information using data from the most correlated channels (Little and Rubin, 2019). This method leverages the inherent similarity between channels to estimate missing data, offering a more informed and context-aware alternative to mean imputation.
@@
$\bullet$ Probabilistic Imputation ($Prob$). Missing channel information is estimated using probabilistic methods (Fekade et al., 2018). By incorporating probabilistic models, it can better handle the complexity and randomness inherent in real-world datasets, making it a robust choice for imputation.
$\bullet$ Deep Interpolation ($AutoEnc$). We leverage the interpolation ability of autoencoders (AutoEnc) to interpolate the complex missing information and then use the interpolated feature as the input for SSL network. We compare against this method to evaluate the effectiveness of deep learning-based interpolation in handling missing data, ensuring that our model’s performance is benchmarked against advanced techniques that utilize neural networks for data reconstruction.
$\bullet$ $SoundSeive$. We apply a state-of-the-art hierarchical interpolation method for acoustic signals, as proposed by SoundSeive (Monjur et al., 2023). Unlike previous time-domain imputation methods, this approach leverages deep learning techniques in the Short-Time Fourier Transform (STFT) domain. By working in the frequency domain, $SoundSeive$ effectively captures and reconstructs complex spectral patterns, leading to a more accurate and robust imputation of missing data. We validate this method to benchmark LOCUS against cutting-edge deep learning approaches, ensuring our model’s performance is compared to the best techniques in both time and frequency domains.
Evaluation Networks. We evaluate each imputation method on two types of networks – (1) pretrained and (2) retrained.
[…]
$\bullet$ Retrained ($ReTrain$). The SSL network is retrained on corrupt data ($Corrupt$) or imputed data ($mean$, $HotDeck$, SoundSieve, and $AutoEnc$) with missing channels for $p$of the time (e.g., MDP=$75$%). Retraining with corrupt data evaluates the model’s performance when it is specifically trained to handle missing data, providing a more robust comparison. It demonstrates whether training the model with data that includes missing channels is sufficient. On the other hand, retraining with imputed data Missing data is imputed data evaluates the effectiveness of these strategies in conjunction with retraining.
5.5. Evaluation Metrics
@@
We evaluate the performance of our sound event localization on the Degree of Arrival Estimation Error ($E_{DoA}$) of the entire dataset. $E_{DoA}$ is a suitable metric for sound source localization because it directly measures the angular difference between the estimated and true locations of sound sources. This angular error provides a clear indication of how accurately the localization system can determine the direction of incoming sounds, which is crucial for applications that rely on precise spatial audio information.
@@ 6. Results
This section begins by evaluating LOCUS on various This section evaluates LOCUS on various localization algorithms, comparing it to a state-of-the-art method to show current limitations. We analyze performance against baselines under varying missing information and conduct an ablation study to assess LOCUS’s components. Finally, we test LOCUS in a different sensor domain to show its generalizability. Section 7 further evaluates LOCUS in real-world environments.
Figure 9. Comparison among different imputation methods with pre-trained models at $MDP=75\%$.
Figure 10. Comparison among different retrained methods at $MDP=75\%$.
Figure 11. Computational complexity and overhead analysis of LOCUS, SELDNet, and SALSA.

### v3

@@ 4.1. InFo: Information-Weighted Focus
Figure 6. Step-by-step workflow of Guided Replacement (GRep) to estimate the full-rank feature $\hat{F}$ from the low-rank feature $\tilde{F}$. Here confidence score $I(\tilde{F})$ acting as weight. When $I_{j,k}$ indicates low confidence in the original $F_{j,k}$, the system relies more on $\hat{F}{j,k}$; conversely, when $F_{j,k}$ appears reliable, it is preserved.
The first component of InFo is the channel confidence branch, $f_{\gamma}(\tilde{F})$, parameterized with $\gamma$. This branch performs average pooling followed by a multi-layer perceptron to extract the channel confidence score. Average pooling aggregates information across all indices, enabling the model to learn more robust features while accurately depicting the overall strength of a feature. The multi-layer perceptron consists of three fully connected layers with batch normalization and rectified linear unit (ReLU) activation after the first two layers. This branch outputs a vector $I_{c}=f_{\gamma}(\tilde{F})$ that measures the global confidence score for each channel.
The second component of InFo is the point-wise confidence branch, $f_{\lambda}(\tilde{F})$, parameterized with $\lambda$. The input $\tilde{F}$ first passes through a convolution layer with a kernel size of $(1\times 1)$ and a channel reduction factor of $r$. It then goes through $n$ dilated convolution layers with a kernel size of $(3\times 3)$, dilation value $d$, batch normalization, and ReLU activation. In our architecture, $n=3$ and $d=2$. The dilated convolution increases the receptive field, enabling us to leverage contextual information. Finally, this contextual information passes through another convolutional layer with a kernel size of $(1\times 1)$ and outputs $I_{s}=f_{\lambda}(\tilde{F})$, which estimates the point-wise confidence score of the input feature.
Next, we broadcast $I_{s}$ to $\mathbb{R}^{d_{\tilde{X}}}$ to match the dimension of $\tilde{F}$ and perform element-wise addition on $I_{c}$ and $I_{s}$ to estimate the information entropy. Finally, the sigmoid function, $\sigma$, maps the results between 0 and 1, where a value closer to 1 indicates higher information availability and vice versa. Therefore, the confidence score is defined as:
@@
4.2. LaFS: Latent Feature Synthesizer
@@
4.3. GRep: Guided Replacement
The Guided Replacement (GRep) module recovers the full-rank feature $\hat{F}$ using Equation 3. Figure 6a depicts this retrieval process. First, we perform element-wise multiplication between the confidence mask $I(\tilde{F})$ and the low-rank feature set $\tilde{F}$ to retrieve the available information in the low-rank feature set $\hat{F}_{\tilde{X}}$.
Next, we find the missing information matrix $\hat{F}_{Z|\tilde{X}}$ from the interpolated feature $\bar{\bar{F}}$. Here, $Z$ represents the missing data in $\tilde{\mathcal{X}}$. As shown in Figure 6b, we take the complement of the confidence mask $I(\tilde{F})$ and multiply it element-wise with $\bar{\bar{F}}$ to find the missing information, disregarding regions (dotted darker green) where errors might have been introduced during interpolation.
@@ 4.3. GRep: Guided Replacement
Dataset Duration $F_{s}$(Hz) No of samples Train-Val-Test
@@
5. Experimental Setup and Implementation Details
This section outlines the datasets, data preparation, implementation, baselines, and evaluation metrics used to assess LOCUS. Our codebase is available at https://github.com/BASHLab/LOCUS.
@@ 5.1. Dataset
We evaluate LOCUS using two distinct multichannel sound source localization datasets, summarized in Table 1. Our evaluation covers both temporal (DCASE dataset, moving sources) and snapshot-based (LargSet, static sources) DoA tasks, demonstrating LOCUS’s generalizability across varied temporal reasoning requirements.
Dataset 1: DCASE. The DCASE2021 Task 4 dataset (Politis et al., 2021) consists of 600 one-minute polyphonic recordings at 24 kHz, captured with a four-channel microphone array. The dataset includes up to four simultaneous sound sources and 12 distinct acoustic events, with Direction of Arrival (DoA) annotations for each source. It also features directional interference and multi-channel ambient noise, with SNR levels ranging from 6 to 30 dB. The dataset is split into 400 training and 200 evaluation samples.
Figure 7. Simulation environment with the microphone array in the center.
Figure 8. Microphone array
@@
Dimension Description
@@
Dataset 2: LargeSet. We created a 50-hour, six-channel dataset with 180,000 one-second recordings at 44.1 kHz across 10 diverse environments. This monophonic dataset avoids overlapping sound sources to isolate the impact of missing channels.
Simulation Details. We simulate a total of ten distinct acoustic environments with PyRoomacoustics (Scheibler et al., 2018) using the configuration parameters outlined in Table 2. Each environment represents a closed rectangular room with rigid boundaries, comprising a ceiling, floor, and four walls, where the material properties and layout may vary across environments to introduce diversity in reverberation and reflection characteristics. The dimensions of each room are defined in meters as $(l\times w\times h)$, representing the length, width, and height, respectively. Figure 8 and figure 8 depict the schematic of the simulation room and the microphone array geometry.
At the geometric center of each room, located at coordinate $\left(\frac{l}{2},\frac{w}{2},\frac{h}{2}\right)$, a spherical microphone array is placed. The array consists of six omnidirectional microphones uniformly distributed along the equator of the sphere. The radius of the microphone array is set to $4.25$ cm, ensuring a compact configuration suitable for near-field sound localization.
To simulate sound sources within each environment, we use the UrbanSound8K dataset (Salamon et al., 2014), to expand the diversity of acoustic content (e.g., engines, alarms, dog barks) and evaluate the generalization of our system beyond typical indoor scenes.. For each simulation instance, a single sound source location is randomly sampled from a continuous uniform distribution within the room volume. Specifically, the source coordinate $(x,y,z)$ is drawn such that $x\in[0,l)$, $y\in[0,w)$, and $z\in[0,h)$, ensuring that the source is located strictly within the room boundaries. This coordinate is recorded as the ground truth for supervised learning tasks such as source localization or direction-of-arrival (DoA) estimation.
@@ 5.2. Data Preparation
Feature Extraction. We extract two key acoustic features: Mel-Frequency Cepstral Coefficients (MFCC) and Generalized Cross-Correlation (GCC) from raw audio using a 1024-point FFT with a 40 ms window and 20 ms hop length at 24 kHz and 44.1 kHz. MFCC captures magnitude-based audio information, while GCC captures time differences and phase shifts between channels, crucial for spatial localization. These features are combined into a matrix, $\tilde{F}$, which is input to the DNN-SSL pipeline.
Missing Channel Perturbation To simulate missing channels, we randomly select $m\leq\frac{n}{2}$ channels from $n$ microphones, where $m$ represents the missing channels. For each Missing Data Percentage (MDP), we replace $p\%$ of the audio with Gaussian noise (mean = 0, std = 1). We also simulate channel failures by randomly dropping entire channels for specific timeframes (100% MDP). These perturbations enable the model to handle real-world scenarios of microphone failure or signal distortion.
@@ 5.3. Implementation Details
Training Pipeline of LOCUS. We train LOCUS jointly with the DNN-based sound source localization (SSL) task. The SSL loss function, $\mathcal{L}_{DT}$, backpropagates through the network and the information branches ($f_{\gamma}$ and $f_{\lambda}$) of InFo, enriching the features used for SSL.
The LaFS branch ($f_{\phi}$) has a separate training objective to estimate the input features. Its loss function, $\mathcal{L}_{DCI}$, is the squared $L_{2}$ distance between the original and reconstructed feature matrices: $\mathcal{L}_{DCI}=||F-\bar{\bar{F}}||^{2}$. This ensures that the reconstructed features closely match the original ones, promoting data consistency. We use two optimizers: one for training InFo and SSL with $\mathcal{L}_{DT}$, and another for training the LaFS branch with $\mathcal{L}_{DCI}$. This dual-optimizer approach ensures optimal performance for both components.
Sound Source Localization Pipeline. We implement two DNN-based SSL solutions for Direction of Arrival (DoA) estimation using multi-channel audio: SELDNet (Shimada et al., 2021; Adavanne et al., 2018) and SALSA (Nguyen et al., 2022), both utilizing two features (Section 5.2).
@@
SALSA is an advanced SSL algorithm with a CRNN model based on the PANN ResNet22 (Kong et al., 2020) architecture. It includes a two-layer Bidirectional GRU and fully connected output layers, adapting to input features by adjusting the number of input channels. During inference, sound classes above the detection threshold are considered active, and their corresponding DOAs are selected.
@@ 5.4. Evaluation Baseline
We evaluate LOCUS by comparing it against four imputation methods and one deep interpolation technique, covering a range of strategies from simple imputation to advanced neural network models. This ensures a thorough comparison of LOCUS ’s effectiveness in handling missing data.
@@
Imputation Methods. We benchmark LOCUS against three signal processing and two deep learning imputation methods, including one designed for acoustic signals. Each method is evaluated on both pre-trained and retrained SSL networks, with retraining done using either corrupt or imputed data.
@@
$\bullet$ Mean Imputation ($Mean$). Replaces missing channel data with the mean of available channels (Allison, 2001), a simple and commonly used baseline method.
@@
$\bullet$ Hot Deck Imputation ($HotDeck$). Fills missing data using the most correlated channels (Little and Rubin, 2019), offering a context-aware alternative to mean imputation.
@@
$\bullet$ Probabilistic Imputation ($Prob$). Estimates missing data using probabilistic models (Fekade et al., 2018), handling complexity and randomness in real-world datasets.
$\bullet$ Deep Interpolation ($AutoEnc$). Uses autoencoders (AutoEnc) to interpolate missing data, comparing its performance in SSL to evaluate the effectiveness of DNN-based interpolation.
$\bullet$ $SoundSeive$. A state-of-the-art hierarchical interpolation method in the STFT domain (Monjur et al., 2023), capturing spectral patterns for robust missing data imputation, providing a benchmark for LOCUS.
Evaluation Networks. We evaluate each method on (1) pretrained and (2) retrained SSL networks.
$\bullet$ Pre-trained ($PreTrain$). The SSL network (e.g., SELDNet) is trained without missing channel data ($MDP=0\%$) and tested on data with varying $p$missing timeframes, assessing performance degradation.
$\bullet$ Retrained ($ReTrain$). The SSL network is retrained with corrupt or imputed data ($mean$, $HotDeck$, SoundSeive, and $AutoEnc$) at different MDP levels, evaluating the robustness of retraining on missing or imputed data.
@@
We evaluate performance using the Degree of Arrival Estimation Error ($E_{DoA}$), which measures the angular difference between the estimated and true sound source locations. $E_{DoA}$ quantifies localization accuracy, providing key insights into system’s ability to determine the direction of incoming sounds, critical for spatial audio applications.
@@ 6. Results
This section begins by evaluating LOCUS on various This section evaluates LOCUS on various localization algorithms, comparing it to a state-of-the-art method to show current limitations. For a fair comparison, we compare against both pretrained and retrained versions of all baseline models under the same missing data conditions using the same training and evaluation splits as LOCUS. This ensures all models are equally exposed to corrupted input during learning. We analyze performance against baselines under varying missing information and conduct an ablation study to assess LOCUS’s components. Finally, we test LOCUS in a different sensor domain to show its generalizability. Section 7 further evaluates LOCUS in real-world environments.
6.1. Evaluation on Different Sound Source Localization Algorithms
Figure 10. Comparison against various imputation methods evaluated using pre-trained models at MDP = 75%. LOCUS achieves the lowest $E_{DoA}$ across both datasets, while the probabilistic imputation method shows the highest degradation, particularly on DCASE, highlighting the robustness of LOCUS under high data loss.
Figure 11. LOCUS achieves the lowest $E_{DoA}$ across both datasets compared to retrained recovery methods with MDP=75%, indicating that simple retraining is insufficient to handle high missingness effectively.
Figure 12. $E_{DoA}$ versus MDPfor PreTrained SELDNet, ReTrained SELDNet, and LOCUS with SELDNet across DCASE and LargeSet. LOCUS maintains significantly lower error across all MDP levels, highlighting its superior robustness to missing data compared to conventional pretraining or retraining approaches.
Table 2. Computational complexity and overhead analysis of LOCUS, SELDNet, and SALSA.